OpenStudy (anonymous):

find the value of lim x->0 {(1-cos4x)/(1-cos6x)}

OpenStudy (anonymous):

\[\frac{\sin^22x}{\sin^23x} =\Large\left(\frac{\frac{2x}{2x}\sin2x}{\frac{3x}{3x}\sin3x}\right)^2 \] 4/9

OpenStudy (anonymous):

gt it tnx, I needed to chek the answers